# Towards foundational LiDAR world models with efficient latent flow matching

### v1

@@ C.2 Semantic Occupancy Forecasting
For semantic occupancy forecasting task, the pretrained CFM model is trained on densified LiDAR frames, denoted by $G_{\theta}^{d}$. The results with numerical value of mIoU and IoU for different pretrained/from-scratch VAE and CFM combinations (NFE=10) at different fine-tuning data fraction are presented in table LABEL:table:_sem_occ_data. It is shown that at 10% of fine-tuning data utilization, our pretrained VAE and CFM with full-parameter fine-tuning provides a 11.17% absolute forecasting performance improvement in mIoU. In general, even the performance gain decays as the amount of fine-tuning data usage increases (which makes sense as it is getting closer and closer to the amount of pretraining data), the effectiveness of our foundational model prevails across all fine-tuning data fractions (at 100%, we still have about 1% performance improvement over the from-scratch model). One key thing to note is that we only fine-tune/train-from-scratch our semantic occupancy forecasting model with half of the available semantic labels for nuScenes (50% of the data used for the same task for other baseline models). This additional splitting of the training set is to prevent the overlap between pretrain and fine-tuning data in terms of geometric structure, as both densified occupancy forecasting and semantic occupancy forecasting models share the same set of scenes and the difference is that semantic occupancy forecasting data is labeled by category. This means that at 10% of fine-tuning data usage, having both VAE and CFM full-parameter fine-tuned based on our foundational model is able to achieve comparable performance with training-from-scratch and OccWorld (<±$\pm 1.5\%$ absolute performance difference in both IoU and mIoU across three-second forecasting horizon) with only 5% of the data it consumes. Please refer to section H for visualizations of the forecasting results.
@@ Appendix D More details on representation alignment
As we mentioned before, during the pre-training phase, we found that it is more appropriate to use new data in fine-tuning the original VAE as opposed to training a VAE from scratch, even though the performance of our vae trained from scratch was excellent. We attribute this to the fact that the pre-trained generative model is in fact more adapted to the data distribution of the original latent, whereas the subspace structure of a latent learned from scratch with fine-tuned data would in fact be very different.
Specifically, for sub-task 3 (nonsemantic to semantic transfer), we cannot directly fine-tune the VAE considering the inconsistent size of the embeddng layers. Instead, we use the approach in Fig. 7: considering that $\bm{o}_{s}$ and $\bm{o}_{d}$ match one-to-one, and the original VAE is fine-tuned on $\bm{o}_{d}$, we can directly constrain the direct outputs of the encoder - mu and sigma. Previous work has tried to use bidirectional KL or EMD distance as a loss, but this does not work in our case: the data for the two modalities are in fact very close, differing only by semantic information.
@@
Therefore we compared the latent from pair-wise $\bm{o}_{d}$ and $\bm{o}_{s}$ in line 1 of Table 13. We can see that both Dual-KL and EMD is very close, only cosine similarity is very different, which actually shown us that the most significant difference in the latent learned from these two VAEs is in fact the direction of the high-dimensional tensors. Then after use cosine similarity to constain the $\bm{o}_{s}$ VAE learning, the result is shown in line 2. With the tensor orientation restriction, we can successfully align the latent of the two VAEs, and the subsequent generative model training demonstrates that our alignment scheme can further improve the results.
@@
For different beam lidar adaptation and outdoor-indoor generalization, while we can directly fine-tune the VAE, we cannot directly compare the similarity of the subspaces (CKAkornblith2019similarityneuralnetworkrepresentations and CKNNAhuh2024platonic require the same or similar samples as inputs to different networks). Our solution is shown in Figure 7. First, we use the original pretrained VAE, denoted by $q(z_{d}|o_{d})$ to get the latent of fine-tuning data without any training(adaption), denoted by $\bm{z}_{old}$. Then use the fine-tuned and training from scratch VAE to obtain the $\bm{z}_{b}$ and $\bm{z}_{a}$. If our hypothesise is correct, the distance between $(z_{a},z_{old})$ should smaller than $(z_{b},z_{old})$.
As data shown in Table D and LABEL:table:task2_vae, we listed the $(z_{a},z_{old})$ and $(z_{b},z_{old})$ in every 2 continuous lines for different fraction of fine-tuning data. We observe that with the increase of data amount used to fine-tuning the VAE, we do can have better result in terms of IOU, even better (when use 100% high beam KITTI data) or very close to the pretrained one. However, what really determines the forecast performance is actually the similarity between the latent space after fine-tuning the VAE and the latent space after only pretrain. Compared to training from scratch, the latent space after fine-tuning the VAE is more similar to the latent produced by using pretrain weight directly, which allows for better use of knowledge from the foundational model. In other words, fine-tuning the VAE is in fact about leaving the original latent space as unaltered as possible while allowing the weights of the original VAE to be adapted (and accurately compressed) to the new sample.
@@ Appendix E Analysis on LoRA fine-tuning with different ranks
To determine the optimal LoRAhu2021loralowrankadaptationlarge configuration for our downstream tasks, we tested the following ranks: 32, 64, and 128 with $\alpha=rank$ and dropout rate of 0.05 on semantic occupancy forecasting subtask with 10% of fine-tuning data and both VAE and CFM components pretrained as explained earlier. We apply LoRA only to the Linear layers, embedding layers, 2D and 3D convolution layers in the CFM architecture. The training is done for 40 epochs for LoRA and full-parameter fine-tuning baseline. The goal of this mini-scale experiment is to find the balance between model performance and parameter efficiency. The result is shown in table 16. First, after sampling the CFM, the model yields suboptimal performance under rank of 32 and 64, with 3.51% and 2.48% absolute performance gap with the full-parameter fine-tuning in terms of mean mIoU across three-second forecasting horizon, respectively. This indicates that the model does not acquire satisfactory level of semantic information from fine-tuning data within the 40 epochs training process which might be caused by lack of depth in the LoRA layers. On the other hand, rank of 128 demonstrates performance that is closest to the full-parameter fine-tuning: 0.25% less in mean IoU and 0.91% less in mean mIoU. This shows that rank of 128 offers enough depth for information to be extracted from the semantic occupancy data and demonstrate comparable performance against the baseline. This rank and $\alpha$ combination is what we present earlier for all downstream tasks under the "LoRA" fine-tuning category.
We did not test any higher or lower rank because: since rank of 32 and 64 both has considerable performance gap compared to the full-parameter fine-tuning, lower rank means less parameters (depth) for LoRA layers which means the model is less likely to learn useful information from the fine-tuning data and consequently, is unlikely to achieve comparable performance with the full-parameter method; on the other hand, rank of 128 already has 10.21 million trainable parameters in total where our CFM model alone only has 11.51 million parameters. This means that under such rank, we are fine-tuning about the same number of parameter as the full-parameter fine-tuning approach. This more-or-less makes experiment the model performance with higher LoRA rank redundant because we will be fine-tuning even more parameters than vanilla fine-tuning and the key advantage of LoRA is to fine-tune the model with optimal performance and reduced trainable parameters. With that said, we tried rank of 32, 64, and 128 and utilizes rank=128 variation for all LoRA fine-tuning completed for our downstream tasks.
@@ F.1 The data efficiency of proposed VAE structure
Follow the official splitting of train/validation set in nuScenes, we test the proposed VAE performance under different fraction of training data. As shown in Table 17, our method only need half of the training data to exceed all of previous data compressor, as we mentioned in the abstract.
@@
F.2 Results of forecasting on pretraining data
In our setting, we use $\bm{o}_{v}$ (Sparse voxels) as pretraining data for sparse-dense beam lidar adaptation, and $\bm{o}_{d}$ (Densified voxels) for the rest of 2 subtasks. Here in Table 18, we show the IOU of these 2 pretraining models which will be used as "foundational" models later.
Table 18: Pretraining performance on different LiDAR representations, without future trajectory.
[…]
1s 2s 3s Avg
$\bm{o}_{v}$ 26.98 21.56 18.26 22.27
$\bm{o}_{d}$ 39.64 29.35 23.73 30.91
Considering the sparsity, it’s naturally more difficult to have a accurate forecast result under $O_{v}$, as the undensified point clouds to lead to a "swimming" effect of points on the surface of objects.
F.3 VAE and forecasting result break down
@@
Now, we present a per-category performance analysis. In Table 20, we compare the average IOU for 3s under each category. Compared to OccWorldzheng2024occworld , our improvement in forecasting performance for medium-sized objects is particularly significant: we improve the forecasts of bicycle, Const. Veh. and Motorcycle by 85.36%, 118.12% and 67.17%, respectively. A broadly performance gain also observed when include future trajectory as condition, the top 2 improvement categories are also small object, relatively 65.18% and 51.2% for Motorcycle and Pedestrian respectively.
F.4 NFE selection for CFM solver during inference
To determine the optimal NFE for our CFM solver during inference, we experiment how the performance and efficiency of our proposed VAE plus CFM architecture for semantic occupancy forecasting vary for different values of NFE. Specifically, for NFE values, we start with NFE=100 and do the decrement of 50, 25, 10, and 5. For performance, as we are performing semantic occupancy forecasting, we infer our model with the corresponding NFE and record mean IoU and mIoU for comparison. For efficiency, we collect two key metrics: Gflops per frame and frame per second for one iteration of sampling that produces 6 frames with 2 Hz of frequency. The results for this experiment is depicted by figure 8.
According to the result, in terms of performance, it is shown model performance does not scale up with NFE. In contrast, we can see that NFE=10 offers the best value of mIoU and second-best value of IoU (0.1% behind the best one at NFE=25). Evaluating efficiency in terms of Gflops and FPS, Gflops per frame increases and FPS decreases with increasing NFE, which suggests the negative correlation between efficiency and NFE. Since NFE=10 gives the best overall performance, we can also observe that it has the second-best efficiency metric overall. For the sampling of our efficiency latent flow matching framework, we want to emphasize model performance and efficiency at the same time to allow better real-world implementation. Thus, based on the result, we choose NFE=10 for all the model inference process.
F.5 Model performance with different fractions of pretraining data
@@
In the previous content we showed that based on the foundational model, the performance boost when we increase the fine-tuning data amount and also the result w/w.o the foundational model. But another question is whether the performance of the model improves with the inclusion of more pretraining data? We designed a simple experiment to answer this question.
@@
We choose 25% of fine-tuning data for 3 different subtasks to simplify the fine-tuning. In Fig. 9, 100% data on the x-axis represents the full pretraining data $O_{v}$. We observe a clear performance improvement (relative improvements of 21.9% , 22.2% and 12.9% for indoor occupancy forecasting, semantic occupancy forecasting and high-beam adaptation tasks) as the amount of pretraining data varies from 50% of $O_{v}$ to 100%.
F.6 CFM performance sensitivity analysis with sampling randomization
The inference process of flow matching is essentially solving the Probability Flow ODE(PF-ODE) as mentioned by previous worksong2023consistency , which is actually deterministic given the sampled Gaussian noise is fixed. Here we also provide a experiment to illustrate this stability. Specifically, using Euler solver, we evaluate the two foundational models (sparse and dense occupancy forecasting) and semantic occupancy forecasting model with 5 random seeds and record the average and standard deviation for IoU and mIoU across the three-second forecasting horizon (on the validation data). We present the results in Table 21, and note that the standard deviation of the performance metrics for all three models is less than 1% of the mean across the entire forecasting horizon. Therefore, we conclude that our CFM model has low performance sensitivity to randomization in the sampling process and the uncertainty around the performance metrics evaluated is minimal.
Table 21: Different CFM models and their average IoU/mIoU (±std) over a three-second horizon, tested with 5 different random seeds.
@@ Appendix G Limitation and future work
Physical consistency. Although our method achieves state-of-the-art performance across different time horizons, the issue of continuous consistency in generated videos remains unresolved. For foreground objects, The current solution only achieves smooth output by performing timing processing within the network structure, but in the generated scenes, especially foreground objects, are still prone to discontinuities. As shown in Table 20, although our work has greatly improved the accuracy of the forecasting foreground objects, there is still no theoretical or design guarantee that the foreground object will be continuous in the output (i.e., it will not suddenly appear or disappear in a few frames).
Multi-agent trajectory forecasting For a safe end-to-end autonomous driving system, it is still very difficult for the current occupancy world model to not only decode the occupancy map after obtaining the future representation, but also forecast the future multi-modal trajectories of surrounding agents. This involves not only the efficient detection of foreground objects and the aforementioned consistency forecasting, but also how to introduce future possibilities into ODE-based sampling methods such as flow matching. We hope that this work will provide an efficient implementation framework for solving this issues in the future and accelerate the progress of subsequent work.

### v2

@@ C.2 Semantic Occupancy Forecasting
For semantic occupancy forecasting task, the pretrained CFM model is trained on densified LiDAR frames, denoted by $G_{\theta}^{d}$. The results with numerical value of mIoU and IoU for different pretrained/from-scratch VAE and CFM combinations (NFE=10) at different fine-tuning data fractions are presented in Table LABEL:table:_sem_occ_data. It is shown that at 10% of fine-tuning data utilization, our pretrained VAE and CFM with full-parameter fine-tuning provides a 11.17% absolute forecasting performance improvement in mIoU. In general, even the performance gain decays as the amount of fine-tuning data usage increases (which makes sense as it is getting closer and closer to the amount of pretraining data), the effectiveness of our foundational model prevails across all fine-tuning data fractions (at 100%, we still have about 1% performance improvement over the from-scratch model). One key thing to note is that we only fine-tune/train-from-scratch our semantic occupancy forecasting model with half of the available semantic labels for nuScenes (50% of the data used for the same task for other baseline models). This additional splitting of the training set is to prevent the overlap between pretrain and fine-tuning data in terms of geometric structure, as both densified occupancy forecasting and semantic occupancy forecasting models share the same set of scenes and the difference is that semantic occupancy forecasting data is labeled by category. This means that at 10% of fine-tuning data usage, having both VAE and CFM full-parameter fine-tuned based on our foundational model is able to achieve comparable performance with training-from-scratch and OccWorld ($\pm 1.5\%$ absolute performance difference in both IoU and mIoU across the three-second forecasting horizon) with only 5% of the data it consumes. Please refer to section H for visualizations of the forecasting results.
@@ Appendix D More details on representation alignment
As we mentioned before, during the pretraining phase, we found that it is more appropriate to use new data in fine-tuning the original VAE as opposed to training a VAE from scratch, even though the performance of our vae trained from scratch was excellent. We attribute this to the fact that the pretrained generative model is in fact more adapted to the data distribution of the original latent, whereas the subspace structure of a latent learned from scratch with fine-tuned data would in fact be very different.
Specifically, for sub-task 3 (nonsemantic-to-semantic transfer), we cannot directly fine-tune the VAE considering the inconsistent size of the embedding layers. Instead, we use the approach in Fig. 8: considering that $\bm{o}_{s}$ and $\bm{o}_{d}$ match one-to-one, and the original VAE is fine-tuned on $\bm{o}_{d}$, we can directly constrain the direct outputs of the encoder: mu and sigma. Previous work has tried to use bidirectional KL or EMD distance as a loss, but this does not work in our case: the data for the two modalities are already very close, differing only by semantic information.
@@
Therefore we compared the latent from pair-wise $\bm{o}_{d}$ and $\bm{o}_{s}$ in line 1 of Table 14. We can see that both Dual-KL and EMD are very close, only cosine similarity is very different, which actually show us that the most significant difference in the latent learned from these two VAEs is in fact the direction of the high-dimensional tensors. Then after using cosine similarity to constrain the $\bm{o}_{s}$ VAE learning, the result is shown in line 2. With the tensor orientation restriction, we can successfully align the latent of the two VAEs, and the subsequent generative model training demonstrates that our alignment scheme can further improve the results.
@@
For different beam LiDAR adaptation and outdoor-indoor generalization, while we can directly fine-tune the VAE, we cannot directly compare the similarity of the subspaces (CKAkornblith2019similarityneuralnetworkrepresentations and CKNNAhuh2024platonic require the same or pair-wise samples as inputs to different networks). Our solution is shown in Figure 8. First, we use the original pretrained VAE, denoted by $q(z_{d}|o_{d})$ to get the latent of fine-tuning data without any training (adaptation), denoted by $\bm{z}_{old}$. Then we use the fine-tuned and training from scratch VAE to obtain the $\bm{z}_{b}$ and $\bm{z}_{a}$. If our hypothesis is correct, the distance between $(z_{a},z_{old})$ should be smaller than $(z_{b},z_{old})$.
As data shown in Table D and LABEL:table:task2_vae, we listed the $(z_{a},z_{old})$ and $(z_{b},z_{old})$ in every 2 continuous lines for different fractions of fine-tuning data. We observe that with the increase of the data amount used to fine-tuning the VAE, we can have better results in terms of IoU, even better (when using 100% high beam KITTI data) or very close to the pretrained one. However, what really determines the forecast performance is actually the similarity between the latent space after fine-tuning the VAE and the latent space after only pretraining. Compared to training from scratch, the latent space after fine-tuning the VAE is more similar to the latent produced by using pretrained weight directly, which allows for better use of the knowledge from the foundational model. In other words, fine-tuning the VAE is in fact about leaving the original latent space as unaltered as possible while allowing the weights of the original VAE to be adapted (and accurately compressed) to the new sample.
@@ Appendix E Analysis on LoRA fine-tuning with different ranks
To determine the optimal LoRAhu2021loralowrankadaptationlarge configuration for our downstream tasks, we tested the following ranks: 32, 64, and 128 with $\alpha=rank$ and a dropout rate of 0.05 on the semantic occupancy forecasting subtask with 10% of fine-tuning data and both VAE and CFM components pretrained as explained earlier. We apply LoRA only to the linear layers, embedding layers, 2D and 3D convolution layers in the CFM architecture. The training is done for 40 epochs for LoRA and full-parameter fine-tuning baseline. The goal of this mini-scale experiment is to find the balance between model performance and parameter efficiency. The result is shown in table 17. First, after sampling the CFM, the model yields suboptimal performance under the rank of 32 and 64, with 3.51% and 2.48% absolute performance gaps with the full-parameter fine-tuning in terms of mean mIoU across three-second forecasting horizon, respectively. This indicates that the model does not acquire a satisfactory level of semantic information from fine-tuning data within the 40 epochs training process which might be caused by lack of depth in the LoRA layers. On the other hand, rank of 128 demonstrates performance that is closest to the full-parameter fine-tuning: 0.25% less in mean IoU and 0.91% less in mean mIoU. This shows that a rank of 128 offers enough depth for information to be extracted from the semantic occupancy data and demonstrates comparable performance against the baseline. This rank and $\alpha$ combination is what we presented earlier for all downstream tasks under the "LoRA" fine-tuning category.
We did not test any higher or lower rank because: since both ranks of 32 and 64 have a considerable performance gap compared to the full-parameter fine-tuning, lower rank means even less parameters (depth) for LoRA layers which makes the model less likely to learn useful information from the fine-tuning data and consequently, is unlikely to achieve comparable performance with the full-parameter method; on the other hand, a rank of 128 already has 10.21 million trainable parameters in total where our CFM model alone only has 11.51 million parameters. This means that under such rank, we are fine-tuning about the same number of parameters as the full-parameter fine-tuning approach.
@@ F.1 The data efficiency of proposed VAE structure
Follow the official split of train/validation set in nuScenes, we test the proposed VAE performance under different fractions of training data. As shown in Table 18, our method only needs half of the training data to exceed all of the previous data compressors, as we mentioned in the abstract.
@@
F.2 VAE and forecasting result break down
@@
Now, we present a per-category performance analysis. In Table 20, we compare the average IoU for 3s under each category. Compared to OccWorldzheng2024occworld , our improvement in forecasting performance for medium-sized objects is particularly significant: we improve the forecasts of bicycle, Const. Veh. and Motorcycle by 85.36%, 118.12% and 67.17%, respectively. A broad performance gain is also observed when including future trajectory as a condition, the top 2 improvement categories are also small objects, relatively 65.18% and 51.2% for Motorcycle and Pedestrian respectively.
F.3 Forecasting performance as a function of the VAE’s compression ratio
We present the reconstruction rates of the VAE architecture at various compression ratios in Table 2. Then a further question arises: which category of information is more susceptible to loss? In Table 21, we show the performance drop for 4 listed reconstruction rates.
Table 21: Proposed VAE per category reconstruction rate break down for different compression ratio
Others
Barrier
Bicycle
Bus
Car
Const. Veh.
Motorcycle
Pedestrian
Traffic Cone
Trailer
Truck
Drive. Surf.
Other Flat
Sidewalk
Terrain
Man-made
Vegetation
x32 98.56 99.38 98.61 99.54 99.68 97.69 99.59 99.46 99.07 98.84 99.21 99.71 99.59 99.21 98.83 97.87 93.66
x192 88.51 97.80 97.45 93.63 96.00 84.19 98.42 97.65 97.68 89.12 92.71 97.27 98.56 94.40 93.86 83.66 68.27
x384 79.56 95.14 96.03 87.75 88.94 75.26 96.34 95.28 95.33 81.89 85.67 93.01 96.61 86.94 86.61 72.52 58.32
x768 66.14 88.74 90.77 79.57 78.89 62.37 90.49 88.07 87.73 69.97 76.63 87.55 90.74 77.29 76.03 62.53 50.56
Drop (x32->x192) 10.05 1.58 1.16 5.91 3.68 13.50 1.17 1.81 1.39 9.72 6.50 2.44 1.03 4.81 4.97 14.21 25.39
Drop (x32->x384) 19.00 4.24 2.58 11.79 10.74 22.43 3.25 4.18 3.74 16.95 13.54 6.70 2.98 12.27 12.22 25.35 35.34
Drop (x32->x768) 32.42 10.64 7.84 19.97 20.79 35.32 9.10 11.39 11.34 28.87 22.58 12.16 8.85 21.92 22.80 35.34 43.10
Interestingly, we notice that the voxel missing in the small and rare foreground objects is much less severe than in the background (vegetation, manmade) and large objects (construction vehicle) during compression ratio increase, while these small targets are more important for safe driving.
Next we show the forecast performance as a function of VAE for different compression rates. First, the results of CFM based on latent with different compression ratios are shown below. For time considerations, all CFMs here are trained for 200 epochs.
Table 22: Forecasting result with different compression ratio VAE
Compression 1 s IoU 1 s mIoU 2 s IoU 2 s mIoU 3 s IoU 3 s mIoU Avg IoU Avg mIoU
x32 38.40 28.96 28.54 18.52 23.14 13.48 30.02 20.31
x192 40.53 33.17 30.37 21.09 24.44 15.64 31.78 23.33
x384 39.68 32.96 29.95 20.91 24.25 14.98 31.29 22.95
x768 34.66 27.10 24.51 17.30 22.61 12.95 27.26 19.11
At the highest reconstruction rate of x32, we find that the network is not fully fitted at 200 epochs, and a longer training time may allow the network to fit better. However, since this is clearly contrary to the ‘sample/training efficiency’ of this paper, we keep the training time the same for fair evaluation. For latents with higher compression multiples, the forecast performance continues to degrade. We believe that this is not only due to the fact that it retains less information on large objects, but also because at larger compression multiples, latents become very sensitive to noise, i.e., the latent space loses smoothness. We find that at multiples of x384 and x768, noise of the order of 1e-4 can crash the network (the magnitude of the latent quantity is approximately between 1e-2 and 1e-3).
F.4 Model performance with different fractions of pretraining data
@@
In the previous section, we showed that based on the foundational model, the performance boost when we increase the fine-tuning data amount and also the result w/w.o the foundational model. But another question is whether the performance of the model improves with the inclusion of more pretraining data? We designed a simple experiment to answer this question.
@@
We choose 25% of fine-tuning data for 3 different subtasks to simplify the fine-tuning. In Fig. 9, 100% data on the x-axis represents the full pretraining data $O_{v}$. We observe a clear performance improvement (relative improvements of 21.9% , 22.2%, and 12.9% for indoor occupancy forecasting, semantic occupancy prediction, and high-beam adaptation tasks) as the amount of pretraining data varies from 50% of $O_{v}$ to 100%.
F.5 CFM performance sensitivity analysis with sampling randomization
The inference process of flow matching is essentially solving the Probability Flow ODE (PF-ODE) as mentioned by previous workssong2023consistency , which is actually deterministic given the sampled Gaussian noise is fixed. Here we designed an experiment which explores the performance change when using different i.i.d. Gaussian noise as input. Specifically, using the Euler solver, we evaluated the semantic occupancy forecasting model with 5 random seeds and recorded the average and standard deviation for IoU and mIoU across the three-second forecasting horizon (on the validation data).
We present the results in Table 23, and note that the standard deviation of the performance metrics for all three models is less than 1% of the mean across the entire forecasting horizon. Therefore, we conclude that our CFM model has low performance sensitivity to randomization in the sampling process and the uncertainty around the performance metrics evaluated is minimal.
Table 23: Different CFM models and their average IoU/mIoU (±std) over a three‑second horizon, tested with 5 different random seeds.
@@ Appendix G Limitation and future work
Physical consistency. Although our method achieves state-of-the-art performance across different time horizons, the issue of continuous consistency in generated videos remains unresolved. For foreground objects, our current solution only achieves smooth output by performing timing processing within the network structure, but in the generated scenes, foreground objects are still prone to discontinuities. As shown in Table 20, although our work has greatly improved the accuracy of the forecasted foreground objects, there is still no theoretical or design guarantee that the foreground object will be continuous in the output (i.e., it will not suddenly appear or disappear in a few frames).
Multi-modality forecasting. For a safe end-to-end autonomous driving system, it is still very difficult for the current occupancy world model to not only decode the occupancy map after obtaining the future representation, but also forecast the future multi-modal trajectories of surrounding agents. As we shown in Table 23, the randomness in flow matching or diffusion (originates from Gaussian noise), under a given condition, the future forecast outputs are in fact quite fixed. This may be solved via more detailed control for foreground and background generation. We hope that this work will provide an efficient implementation framework for solving this issue in the future and accelerate the progress of subsequent work.